OpenStudy (anonymous):

Solve x-y=3 and x+y=5

OpenStudy (jdoe0001):

well.. have you covered system of equations yet?

sammixboo (sammixboo):

Well, what two numbers can be subtracted and get the difference of 3, and also be added together to get the sum of 5?

TheSmartOne (thesmartone):

You can always try plugging in numbers :)

OpenStudy (anonymous):

I need to know where I need to add y-3 to both side of equation

OpenStudy (anonymous):

You can solve this by solving for x in terms of y and then substituting that in order to solve for y. Let's solve for x. x+y=5 -y -y ~ Subtract y from both sides. x=5-y ~ Now plug this in to the first equation. (5-y)-y=3 ~ Simplify 5-y-y=3 5-2y=3 ~ Subtract 5 from both sides. -2y=3-5 -2y=-2 y=1 I think you can do the rest! :)